OpenStudy (anonymous):

I WILL FAN !!!!!!! AND MEDAL!!!! The coordinates below represent two linear equations. How many solutions does this system of equations have? Line 1 x y –1 6 0 4 Line 2 x y 0 7 1 5 A. 0 B. exactly 1 C. exactly 2 D. infinitely many

OpenStudy (texaschic101):

LINE 1. find slope of the line using slope formula and points slope (m) = (y2 - y1) / (x2 - x1) (-1,6)...x1 = -1 and y1 = 6 (0,4)...x2 = 0 and y2 = 4 now we sub slope(m) = (4 - 6) / (0 - 1) slope(m) = -2/-1 slope = 2 now we will find the equation of the line by using y = mx + b and 1 set of points...it does not matter which set...(-1,6)...x = -1 and y = 6 slope (m) = 2 now we sub and find b (the y intercept) y = mx + b 6 = 2(-1) + b 6 = -2 + b 6 + 2 = b 8 = b so this line is : y = 2x + 8 ---------------------- LINE 2 do exactly the same thing slope (m) = (5 - 7) / (1 - 0) slope(m) = -2/1 or just -2 y = mx + b slope(m) = -2 (0,7)...x = 0 and y = 7 now sub 7 = -2(0) + b 7 = 0 + b 7 = b so this line is : y = -2x + 7 --------------------- so our 2 lines are : y = 2x + 8 and y = -2x + 7 In y = mx + b form, the m stands for the slope and the b stands for the y intercept. We will now compare the slopes and the y intercepts. Line 1 : slope = 2 and y intercept = 8 Line 2 : slope = -2 and y intercept = 7 Learn this : If the slopes are the same and the y intercepts the same, it will have infinite solutions. If the slopes are different, then there is 1 solution If the slopes are the same and the y intercepts different, it will have 0 solutions So knowing this, what do you think the answer is ?